OpenStudy (jack1):

k here's a weird one: i needs to convert these ones (or maybe just show me how to do the first one if i pick it up like lightning ;) ) from decimal to binary... but it's literally a decimal...? 0.98 0.347 0.9028

OpenStudy (jack1):

hay again ;D

ganeshie8 (ganeshie8):

\(\large 0.98 \times 2 = \color{red}{1}.96\) \(\large 0.96 \times 2 = \color{red}{1}.92\) \(\large 0.92 \times 2 = \color{red}{1}.84\) ...

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=0.98++to+binary

OpenStudy (johnweldon1993):

Oh oh oh oh wait! lol I remember this now What you do...is you continuously multiply by 2 to break it down.. .98 x 2 = 1.96 that whole number will be your first binary digit to the right of the decimal point... now throw out that 1...so we have .96 x 2 = 1.92 again we have a 1....so so far we have .11 for our binary numbers and again we keep going and going .92 x 2 = 1.84 so we have another 1 (can you tell this will go on for a while? we have to keep going until we get down to the 0's but it looks like this will be an infinite binary fraction

OpenStudy (jack1):

awesome, thnks @johnweldon1993 and @ganeshie8 , i just got back on as OS booted me off for an hour, but that seems pretty easy to follow